OpenStudy (anonymous):

Is arcsin(3/3) equals pi/2 why?

OpenStudy (freckles):

3/3=1 arcsin(1)=x where x is between -pi/2 and pi/2 (inclusive) taking sin( ) of both sides gives 1=sin(x) the only solution to this in [-pi/2,pi/2] is x=pi/2 since sin(pi/2)=1 so arcsin(1)=pi/2